# Performances in solving the Bethe-Salpeter equation with the Yambo code

### v1

@@ 2.2 The iterative solver
MPI parallelization is based on standard row-wise one-dimensional partitioning of matrices and vectors. This is the scheme used by PETSc, which is the library on which SLEPc is built. In the case of GPU simulations, the underlying local computations are offloaded to the NVIDIA GPU using CUDA (or HIP for AMD GPUs). The most expensive step of the algorithm are the matrix-vector products that generate the new vectors of the basis. These vectors need to be generated one by one, as they have to be orthogonalized against the previous ones. Additionally, although the solver is also optimized in the case of a sparse matrix, the problems that arise from applications are typically dense, resulting in an expensive matrix-vector product. The details of the SLEPc solvers for the Bethe–Salpeter eigenproblem are available in [35].
@@ 3.1 Details on the performed simulations
The characteristics of two clusters used for CPU-only and GPU simulations are reported in Table 1. We built in single precision (32-bit float numbers for real and imaginary parts) matrices of Eq. (1) for a single layer of chromium triiodide CrI3. The set of BSE-like matrices considered in the present report are generated by replacing the demanding BSE kernel with the cheaper TD-DFT kernel. We used an energy cut-off of 4 Ry for the eh-exchange $v$, and 10 Ry for the $f_{xc}$ kernel. These parameters determine the pre-factor for the building time of the BSE kernel which we do not analyze here. The size of the BSE matrix is defined by the number of bands used. We considered the three cases reported in Table 2. Here, we focus on the performance of the two solvers described in Sect. 2. For direct diagonalization, ScaLAPACK and ELPA libraries (solver type 2 on CPU, type 1 on GPU) were used with a block size of $64\times 64$ for the block-cyclic distribution. For the iterative solver, the SLEPc library was used, with either explicit (e) or shell (s) matrix formats.
Table 2: Relationship between memory, matrix size and BSE bands. Memory estimated assuming 64-bit single precision floating point complex numbers.
@@
[…]
@@
  • Strong scaling plots, i.e. time ($t_{i}$) vs number of MPI tasks ($n_{i}$) at fixed matrix size $N_{i}$.
@@
  • Memory scaling plots, i.e. host memory ($M_{i}$) vs number of MPI tasks ($n_{i}$) at fixed matrix size ($N_{i}$).
@@
  • Efficiency plots, i.e. efficiency ($e_{i}(j)$) vs number of MPI tasks ($n_{i}$) at fixed matrix size. For the CPU case $n_{j}=n_{\text{CPU}}=4$, while for the GPU case $n_{j}=n_{\text{GPU}}=2$.
@@
For an efficiency plot, when a reference value is not available, we estimate it using the exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, where $t_{1}$ represents the time required for a run on a single MPI task, and $p$ is a constant that characterizes how the simulation time scales with the number of processors. For perfect scaling, $p=-1$, i.e., doubling the number of MPI tasks should halve the simulation time. Any value $p>-1$ indicates inefficiencies in parallelization, with a time reduction lower than expected. We will refer to the fitted parameter $p$ as the parallelization factor.
@@
For both CPU and GPU computations, we measured memory usage using the Linux time command, which reports peak memory of each MPI task. For some of the runs we also monitored the peak memory of the run reported by the PETSc internal memory tracking tools. To analyze the memory use, we assume that the allocated host memory is due to arrays/matrices which are either perfectly distributed or fully duplicated. We define the distributed memory as the intercept, and the duplicated memory as the slope of the linear fit of the allocated memory vs number of MPI tasks. Moreover, since in some simulations we observe a sudden memory jump from the serial to the parallel simulations, we also define the parallel–overhead memory as the difference between the fit intercept and the memory allocated in the serial run. We attribute this additional memory increase in parallel runs to the allocation of extra variables not present during serial execution. Memory extracted by the time command and the PETSc tracker agree well, with a parallel overhead detected only by the time command.  [^3] The PETSc data are slightly smaller than the data extracted with the time command in parallel runs. This possibly happens because PETSc tracks the peak memory of the whole run, while time tracks the peak memory of each MPI task, which can be reached at different moments during the simulations.
@@
[…]
Table 3: Exponential fit of time vs number of CPUs, $t_{i}=t_{1}\,n_{i}^{p}$. Where $p=-1$ is the ideal parallelization case. (e) - explicit matrix type, (s) - shell matrix type.
@@
Fig. 1 shows that the ELPA library outperforms ScaLAPACK library. This is also reflected in the p values reported in Table 3, which are closer to -1 for ELPA. In fact, among all the studied solvers, ELPA demonstrates one of the best efficiencies and scaling behaviors. As the matrix size increases, ELPA’s efficiency will plateau depending on the size of matrix and type of matrix, with larger matrices benefiting more from ELPA’s optimization, whereas for the ScaLAPACK the plateau always happens around the value of efficiency in the range 0.4 - 0.5.
Figure 1: Execution time, memory usage and efficiency plots by different solvers and different matrices. CPU case.
SLEPc generally runs about an order of magnitude faster than its full diagonalization counterparts, here we ask to extract 100 eigenvalues. The efficiency of the SLEPc solver is in general worse than the diagonalization case. The shell matrix algorithm, in particular, performs very poorly, with a total time which starts to grow after a critical number of MPI tasks is reached. This is however, in part, expected. Since in the shell approach the algorithm uses a matrix distribution which is optimized for building of the kernel. The explicit approach is instead more efficient, although not as efficient as full diagonalization.
As shown in Fig. 1, efficiency drops from 1 to about 0.4 when using up to 16 cores. Beyond that point, efficiency plateaus for the explicit case (while continuing to decline for the shell case). Additionally, for SLEPc, neither the type nor the size of the matrix significantly affects efficiency, which consistently plateaus around 0.4.
Duplicated memory remains consistently around 0.4–0.5 GB per CPU across all cases, likely due to internal Yambo operations and is independent of the solver. Its impact decreases with matrix size. In contrast, distributed memory depends on both solver and matrix size. Direct diagonalization shows high memory usage – 5–6 times larger than the calculated value for resonant case (Table 2); and coupling case uses about 3-4 times more memory than the resonant case. SLEPc requires much less memory, enabling single MPI runs, unlike full diagonalization. In the explicit case, memory usage increases notably when transitioning from serial to parallel runs which we attribute to parallel overhead (see Table 4). This effect is likely due to additional variable allocations. Reducing this overhead could enhance SLEPc’s efficiency.
@@
Figure 2: Execution time, peak memory, and efficiency plots by different solvers on matrices of 3 different sizes. GPU case. For SLEPc, coupling matrix, data extrapolated assuming $\mathrm{time}=\frac{4973}{1.7^{\log_{2}(\mathrm{nr\_gpu})}}$. Only host memory is reported here.
Table 5: Exponential fit of time vs number of GPUs, $t_{i}=t_{1}\,n_{i}^{p}$. $p=-1$ is the ideal parallelization case.
@@
Table 6: Distributed (Distr.) and duplicated (Dupl.), and parallel overhead (Par.) memory for the GPU case. All units in [GB].
@@
ELPA res 5.02 2.08 -0.16 64 2.00 -0.17 526 2.00
ELPA cpl 16.51 2.03 -0.38 208 1.98 - 1720 1.85
  SLEPc res 4.13 2.11 4.36 72 2.45 63.92 861 1.95
SLEPc cpl 6.32 2.15 5.68 102 3.45 71.89 1172 2.26
[…]
Medium-sized matrices are handled more efficiently. Their parallelization coefficient is higher (closer to the values observed in the CPU case, see Tables 3 and 5). For these matrices, a performance plateau is reached around 32 GPUs for the resonant case and around 64 GPUs for the coupling case. Up to the plateau, the scaling is very good: doubling the number of GPUs roughly halves the execution time, though the plateau is reached relatively quickly.
For the big matrices the scaling is even better. This is evident from the efficiency plots in Fig. 2, and from the parallelization factors in Table 5. We do not observe any plateau up to the maximum number of GPU used, e.g. 64 (128) for the resonant (coupling) case. Overall we can conclude that, the larger the matrices, the better the scaling and the effciency.
Similar to the CPU case, SLEPc executes an order or two of magnitude faster than ELPA. For the SLEPc solver oscillations in the execution time have been observed. No consistent pattern was found, but data suggest that numbers of GPU equal to powers of $2$ perform better.
Memory usage analysis shows that duplicated memory remains consistently around 2 GB per GPU, which is 4 times higher than for the CPU case. The distributed memory is very close to the one obtained in the CPU case, see Table 4. On Leonardo Booster, per each GPU there is 64GB of memory, and for the cases of medium and big matrices, this imposes a minimal amount of GPUs to perform the simulations. In comparison with the CPU case, on GPUs the SLEPc solver shows similar or even higher memory consumption than the ELPA solver. This is due to the large parallel overhead that we are presently investigating.
For the GPU implementation, the parallel overhead is of a similar order of magnitude to that observed in the CPU case. Specifically, for SLEPc, the parallel overhead is large and positive, indicating the allocation of additional variables for parallel runs. Conversely, for ELPA, the parallel overhead is negative, suggesting more efficient parallelization schemes.

### v2

@@ 2.2.2 MPI Distribution and GPU Porting in SLEPc
The most expensive step of the algorithm are the matrix-vector products that generate the new vectors of the basis. These vectors need to be generated one by one, as they have to be orthogonalized against the previous ones. Additionally, although the solver is also optimized in the case of a sparse matrix, the problems that arise from applications are typically dense, resulting in an expensive matrix-vector product.
@@ 3.1 Details on the Performed Simulations
Table 1: Characteristics of the clusters used to perform the simulations.
The characteristics of two clusters used for CPU-only and GPU simulations are reported in Table 1. We constructed matrices from Eq. (1) in single precision (using 32-bit float numbers for both real and imaginary parts) for the case $A=R$ for a single layer of chromium triiodide CrI3. The set of BSE-like matrices considered in the present report are generated by replacing the demanding BSE kernel with the cheaper TD-DFT kernel [1, 43], as implemented in the Yambo code. Specifically, the matrix blocks are defined as follows:
$R_{i,i^{\prime}}$ $=$ $\Delta\epsilon_{i}\delta_{i,i^{\prime}}+v_{i,i^{\prime}}+f^{xc}_{i,i^{\prime}},$ (8)
$C_{i,j^{\prime}}$ $=$ $v_{i,j^{\prime}}+f^{xc}_{i,j^{\prime}}.$ (9)
Here, $i=\{nm\mathbf{k}\}$ is a transition index. In the resonant block both $i$ and $i^{\prime}$ refer to positive energy transitions from occupied to empty states, $n\mathbf{k}\rightarrow m\mathbf{k}$, while in the coupling block $j^{\prime}$ refers to negative energy transitions, $m\mathbf{k}\rightarrow n\mathbf{k}$, and $\Delta\epsilon_{i}=\epsilon_{m\mathbf{k}}-\epsilon_{n\mathbf{k}}$ are energy differences. We used an energy cut-off of 4 Ry for the eh-exchange $v$, and 10 Ry for the $f_{xc}$ kernel. These parameters determine the prefactor for the building time of the BSE kernel which we do not analyze here. The size of the BSE matrix is defined by the number of bands used. We considered the three cases reported in Table 2. Here we focus on the performance of the two solvers described in Sect. 2. For direct diagonalization, ScaLAPACK and ELPA libraries (solver type 2 on CPU, type 1 on GPU) were used with a block size of $64\times 64$ for the block-cyclic distribution. For the iterative solver, the SLEPc library was used, with either explicit (e) or shell (s) matrix formats. In our experience, when the eigenvectors are extracted, if more than 5% of the spectrum is required, direct diagonalization routines are recommended. For smaller fractions of the spectrum, the SLEPc iterative solvers are preferable. Here we fix the number of extracted eigenvectors to 100. In the parallel runs this ranges from 1% (for the smallest matrices), to 0.1% (for the largest considered matrix) of the spectrum.
@@
Table 2: Relationship between memory, matrix size and BSE bands for resonant (res.) and coupling (cpl.) matrices. Memory estimated assuming 64-bit single precision floating point complex numbers.
In order to analyze the performances of the simulations, we define the following quantities. $N_{i}$: number of rows in the resonant block of the matrix for the run $i$. $n_{i}$: number of MPI tasks for the run $i$. $t_{i}$: total simulation time for the solver step for the run $i$. $e_{i}(j)=\frac{n_{j}t_{j}}{n_{i}t_{i}}$: efficiency of the run $i$ with $n_{i}$ MPI tasks w.r.t. the run $j$ with $n_{j}$ MPI tasks. $M_{i}$: the peak memory or maximum resident memory on the host RAM, obtained by summing over the peak memory of all MPI tasks, for the run $i$. We will show, both for CPU and GPUs runs:
  • Time complexity plots, i.e., total simulation time ($t_{i}$) on a single CPU vs matrix size ($N_{i}$).
@@
  • Strong scaling plots, i.e., time ($t_{i}$) vs number of MPI tasks ($n_{i}$) at fixed matrix size $N_{i}$.
@@
  • Memory scaling plots, i.e., host memory ($M_{i}$) vs number of MPI tasks ($n_{i}$) at fixed matrix size ($N_{i}$).
@@
  • Efficiency plots, i.e., efficiency ($e_{i}(j)$) vs number of MPI tasks ($n_{i}$) at fixed matrix size. For the CPU case $n_{j}=n_{\text{CPU}}=4$, while for the GPU case $n_{j}=n_{\text{GPU}}=2$.
@@
For an efficiency plot, when a reference value is not available, we estimate it using the exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, where $t_{1}$ represents the time required for a run on a single MPI task, and $p$ is a constant that characterizes how the simulation time scales with the number of MPI tasks  [44].
For perfect scaling, $p=-1$, i.e., doubling the number of MPI tasks should halve the simulation time. Any value $p>-1$ indicates inefficiencies in parallelization, with a time reduction lower than expected. We will refer to the fitted parameter $p$ as the parallelization factor.
@@
For both CPU and GPU computations, we measured memory usage using the Linux time command, which reports peak memory of each MPI task. For some of the runs we also monitored the peak memory of the run reported by the PETSc internal memory tracking tools. To analyze the memory use, we assume that the allocated host memory is due to arrays/matrices which are either perfectly distributed or fully duplicated. We define the distributed memory as the intercept, and the duplicated memory as the slope of the linear fit of the allocated memory vs number of MPI tasks. Moreover, since in some simulations we observe a sudden memory jump from the serial to the parallel simulations, we also define the parallel–overhead memory as the difference between the fit intercept and the memory allocated in the serial run. We attribute this additional memory increase in parallel runs to the allocation of extra variables not present during serial execution. Memory extracted by the time command and the PETSc tracker agree well, with a parallel overhead memory detected only by the time command. [45]
3.2 Comparison of non-Hermitian vs pseudo-Hermitian Algorithms
Figure 1: Time complexity (panels (a)–(c)) and memory complexity (panels (d)–(f)) for solving the BSE eigenvalue problem using a single CPU. Results are shown for: (a,d) the Hermitian case, (b,e) the coupling case with a non-Hermitian algorithm, and (c,f) the coupling case with a pseudo-Hermitian algorithm. In all cases, $N$ denotes the size of the resonant block. Simulations were carried out on the ISM cluster.
Before analyzing the CPU and GPU scaling of the simulations, we first analyze how the computational time scales with matrix size (i.e., time complexity) using a single CPU. We compare the standard algorithms for Hermitian (H) matrices (corresponding to resonant case), standard algorithms for non-Hermitian matrices (NH), and novel algorithms that exploit the pseudo-Hermitian (PH) structure of the coupling matrix. Here we consider three different solvers: (i) full diagonalization which is based on (Sca)LAPACK [46] (see Sect. 2.1.1), (ii) Krylov-Schur iterative solver based on the SLEPc library (see Sect. 2.2), and (iii) Haydock iterative solver. The Haydock solver performance is reported for reference as an example of an iterative algorithm implemented in Yambo, which already exploits the pseudo-Hermitian structure of the matrix [41].
Fig. 1 presents time complexity plots, and Table 3 reports the corresponding slope and intercept values, $log(t)=A\cdot log(N)+B$. By moving from a logarithmic representation to the direct time versus matrix size relation, the slope A corresponds to the scaling exponent of the algorithm with respect to matrix size. Lower values of A indicate better scaling. The intercept B represents the computational prefactor, which is also more favorable when smaller. The relation is expressed as $t=B\,N^{A}$. The expected scaling is $\mathcal{O}(N^{2})$ for SLEPc and $\mathcal{O}(N^{3})$ for exact diagonalization. Fig. 1 shows that full diagonalization is the slowest algorithm, while Haydock is the fastest, and SLEPc is intermediate. The performance differences between SLEPc and Haydock arise from the fact that the Haydock algorithm only extracts the optical spectrum, while the SLEPc algorithm also extracts a few eigenvalues and eigenvectors [11].
Method Kind Slope(A) Intercept(B) Kind Slope(A) Intercept(B)
GFortran Compiler (ISM-CNR Cluster)
LAPACK H 2.61 -7.91 NH 3.05 -7.71
PH 2.89 -8.00
ELPA H 2.71 -8.17 PH 2.79 -7.99
SLEPc Expl. H 2.12 -6.33 NH 2.45 -6.30
PH 2.35 -6.64
Haydock H 0.95 -2.77 PH 1.66 -5.35
Intel Compiler (University of Valencia Cluster)
LAPACK H 2.56 -7.81 NH 3.04 -8.12
PH 2.93 -8.29
SLEPc Expl. H 2.19 -6.63 NH 2.85 -7.63
PH 2.36 -6.74
Haydock H 1.37 -4.32 PH 1.88 -5.80
Table 3: Slope (A) and intercept (B) parameters from linear regressions of the time-complexity data for the resonant (H) and coupling cases (non-Hermitian (NH) and pseudo-Hermitian (PH)). The upper section reports results obtained on the ISM cluster using the GFortran compiler; these are the data shown in Fig. 1. The lower section presents reproduced benchmarks performed on the University of Valencia cluster (UVC) using Intel compilers and MKL libraries.
By comparing panels (b) and (c) of Fig. 1, we observe a significant reduction in the time to solution when switching from the NH to the PH algorithm, for both exact diagonalization and the SLEPc based iterative solver. For the SLEPc case a five- to six-fold improvement in speed is obtained. As shown in Table 3, the improvements are due to the better scaling exponential (i.e., slope). In the NH implementation, the scaling behavior was similar to that of full diagonalization algorithms, and the performance advantage over diagonalization was mainly due to a smaller prefactor (i.e., a lower intercept in the scaling fit). For the diagonalization case the speed-up from the NH to the PH case is also roughly a factor five, while the improvement is due to both a slightly better scaling and prefactor [47].
In Table 3 we also put the extracted slope and intercept from simulations done on a different machine, using Intel compiler with MKL libraries, in order to verify how much these numbers are machine dependent. For a given algorithm, the slope A is expected to be independent of the computational configuration (such as system architecture, compiler, and libraries), as it reflects the intrinsic scaling behavior of the algorithm. In contrast, the intercept B can be seen as the product of two factors: $B_{1}$, which depends on the configuration, and $B_{2}$, which is algorithm-dependent. The results confirm that the extracted slope is rather similar between the two machines. In this case the extracted value is $B_{1}^{ISM}/B_{1}^{UVC}\approx 0.93$, which suggests that the two clusters have rather similar performances.
Panels (d), (e), and (f) in Fig. 1 and Table 4 illustrate the memory complexity of the studied algorithms, that is, how memory usage increases with matrix size. The following observations can be made: (i) iterative algorithms consume less memory than diagonalization-based algorithms; (ii) memory complexity is similar between LAPACK and ELPA, and between SLEPc and Haydock; and (iii) the PH implementations result in improved memory efficiency vs the NH implementation for the SLEPc algorithms, while causing a slight increase in memory usage for the LAPACK-based approach [48]. Panels (d) and (f) show that the ELPA and LAPACK memory fits overlap almost perfectly. In contrast, a small difference is observed between Haydock and SLEPc, with the Haydock algorithm being slightly more memory-efficient. Overall the memory increase as a function of the matrix size behaves as expected.
Method kind Slope(A) Intercept(B) kind Slope(A) Intercept(B)
LAPACK H 1.31 -4.64 NH 1.52 -5.16
PH 1.68 -5.56
SLEPc Expl. H 0.72 -2.76 NH 1.23 -4.37
PH 1.03 -3.37
ELPA H 1.35 -4.77 PH 1.68 -5.55
Haydock H 0.67 -2.59 PH 0.81 -3.02
Table 4: Slope and intercept values for the memory complexity plots shown in Fig. 1(d)–(f), for the resonant (res.) and coupling (cpl.) cases. H – Hermitian, NH – non-Hermitian, PH – pseudo-Hermitian. Expl. - explicit matrix type.
Figure 2: Comparison of time complexity for the studied algorithms, with the x-axis representing the actual matrix size: N for the resonant case and 2N for the coupling case.
Fig. 2 shows the time complexity plots rescaled to reflect the actual size of the coupling matrix. As already observed, iterative solvers are orders faster than diagonalization-based approaches. When the full matrix size is considered, instead of only the resonant block size, it can be seen that the two algorithms (H vs PH) behave rather similar performances. The ELPA implementation is the only one where the PH is clearly faster than the H for all matrix sizes considered here. For the iterative solvers, the Hermitian version exhibits a more favorable slope compared to the pseudo-Hermitian one, and this difference becomes more significant as the matrix size increases.
In summary, this subsection demonstrates that, both for SLEPc and diagonalization, the new PH algorithm greatly improves the time to solution at fixed matrix size compared to the NH algorithm, making the solution of BSE with coupling nearly as efficient as the resonant approximation. Memory usage is also improved for SLEPc, while it remains almost unchanged in the diagonalization case.
@@ 3.2 Comparison of non-Hermitian vs pseudo-Hermitian Algorithms
Figure 3: Execution time (panels (a) and (b)), efficiency (panels (c) and (d)), and memory usage (panels (e) and (f)) for different solvers applied to the resonant and coupling cases of the BSE eigenvalue problem. CPU case.
Now we analyze MPI scaling of the solvers on CPU. We consider strong scaling with matrices of size $N=10^{4}$ and $N=3\cdot 10^{4}$. For the coupling case, this corresponds to matrices of size $2N$. In Fig. 3 time and memory scaling plots are presented, together with efficiency plots. The exponential fit of the data is reported in Table 5. Finally, the extracted parameters from the memory fit are reported in Table 6. In this and next subsection, the term scaling is referring to scaling with respect to number of CPUs or GPUs.
@@
Table 5: Exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, of time vs number of CPUs for resonant (res.) and coupling (cpl.) matrices. Where $p=-1$ is the ideal parallelization case. (e) – explicit matrix type, (s) – shell matrix type.
@@
Fig. 3 shows the execution time of iterative and diagonalizational algorithms (see panels (a) and (b)). However, the efficiency of the SLEPc solver is in general worse than the diagonalization case. As shown in Fig. 3, panels (c) and (d), efficiency drops from 1 to about 0.4 when using up to 16 cores. Beyond that point, efficiency plateaus for the explicit case, while continuing to decline for the shell case. The efficiency of the diagonalization solvers instead ranges in between 0.5 and 0.9. The ELPA library consistently shows the highest efficiency and scaling behaviors (see panels (c) and (d)).
Analyzing in detail different flavours of the iterative and diagonalization solvers we observe the following. For the SLEPc case, the shell matrix algorithm performs very poorly, with a total time which starts to grow when a critical number of MPI tasks is reached. This behaviour is somewhat expected as the shell approach uses a matrix distribution which is optimized for building the kernel. Instead, the explicit approach is more efficient, although not as efficient as full diagonalization. For the diagonalization case, the ELPA library outperforms ScaLAPACK, likely due to its more efficient parallelization. This performance difference can be seen in Fig. 3, panel (a). As the matrix size increases, ELPA’s efficiency will plateau in the range of 0.6 – 0.8, depending on the matrix type and size, with larger matrices benefiting more from ELPA’s optimization, whereas for ScaLAPACK, the plateau the plateau consistently occurs around 0.4 – 0.5. In the coupling case, panel (b), ELPA shows an even greater performance advantage. This is most likely due to its more optimized implementation of the pseudo-Hermitian solver, compared to ScaLAPACK. Interestingly, in some cases, ELPA’s efficiency in the pseudo-Hermitian case is higher than in the Hermitian case. These differences are also reflected in the p values reported in Table 5, where ELPA values are closer to the ideal p = $-1$, typically around $-0.8$ to $-0.9$, followed by ScaLAPACK with values in between $-0.7$ and $-0.8$, and finally the SLEPc explicit solver reaches values in between $-0.6$ and $-0.7$. The SLEPc shell solver, again, shows the worst performance in this comparison.
Duplicated memory remains consistently around 0.4–0.5 GB per CPU across all cases, likely due to internal Yambo operations and is independent of the solver, panels (e) and (f). Its impact decreases with matrix size. In contrast, distributed memory depends on both solver and matrix size. Direct diagonalization shows high memory usage approximately 5–6 times larger than the calculated value for resonant case (see Table 2); and coupling case uses about 3–4 times more memory than the resonant case. SLEPc requires much less memory, enabling single MPI runs, unlike full diagonalization. In the explicit case, memory usage increases notably when transitioning from serial to parallel runs which we attribute to parallel overhead memory (see Table 6). This effect is likely due to additional variable allocations. Reducing this overhead could enhance SLEPc’s efficiency.
@@
We now analyze performance and MPI scaling of the solvers on GPU. Here we consider matrices of size $N=10^{4}$, $N=3\cdot 10^{4}$, and $N=10^{5}$, which we refer to as small, medium, and large matrices, respectively. To this end, we focus on the ELPA and SLEPc explicit solvers, since the ScaLAPACK library is not GPU ported, and the SLEPc shell solver is not efficiently scaling. Similar to the CPU case, SLEPc executes an order or two of magnitude faster than ELPA, Fig.  4, panels (a) and (b).
Figure 4: Execution time (panels (a) and (b)), efficiency (panels (c) and (d)), and memory usage (panels (e) and (f)) for different solvers applied to the resonant and coupling cases of the BSE eigenvalue problem. GPU case. Only host memory is reported here.
@@
Table 7: Exponential fit of time vs number of GPUs, $t_{i}=t_{1}\,n_{i}^{p}$. The case $p=-1$ represents the ideal parallelization. The results marked with may be influenced by the limited number of data points in the low GPU count range.
Matrices of size $N\approx 10^{4}$ are too small for the available resources on GPU, and performances of the solvers are dictated by extra operations which prevents them from scaling efficiently. For the resonant case, both ELPA and SLEPc complete a calculation in under one minute on a single GPU (compared to 1–10 minutes on a single CPU). Consequently, performance quickly reaches a plateau: adding more GPUs (beyond 16) yields little to no improvement in execution time. A similar trend is observed for the coupling case, where the plateau begins at 32 GPUs (see Fig. 4). Table 7 shows that for a matrix of size $N\approx 10^{4}$, the parallelization coefficient p is significantly lower than that for the CPU case. Medium-sized matrices are handled more efficiently. Their parallelization coefficient is higher, and is closer to the values observed in the CPU case (see Tables 5 and 7). For these matrices, a performance plateau is reached around 32 GPUs for the resonant case and around 64 GPUs for the coupling case. Up to the plateau, the scaling is very good: doubling the number of GPUs roughly halves the execution time, though the plateau is reached relatively quickly. For the big matrices the scaling is even better. This is evident from the efficiency plots in Fig. 4, panels (c) and (d), and from the parallelization factors in Table 7. We do not observe any plateau up to the maximum number of GPU used, i.e., 64 (128) for the resonant (coupling) case. Overall we can conclude that, the larger the matrices, the better the scaling and the efficiency.
For the ELPA solver, similarly to the CPU-scaling case, by comparing the resonant and coupling cases, Table 7, we observe that it performs significantly better in the coupling case.
For SLEPc, no significant difference was observed between the resonant and coupling cases. Moreover, oscillations in the execution time can be observed suggesting that numbers of GPU equal to powers of $2$ perform better.
@@
ELPA res. 5.02 2.08 -0.16 64 2.00 -0.17 526 2.00
ELPA cpl. 16.51 2.03 -0.38 208 1.98 - 1720 1.85
  SLEPc res. 4.13 2.11 4.36 72 2.45 63.92 861 1.95
SLEPc cpl. 6.32 2.15 5.68 102 3.45 71.89 1172 2.26
Table 8: Distributed (Distr.), duplicated (Dupl.), and parallel overhead (Par.) memory for the GPU case. All units in [GB]. For the largest matrix size, the parallel overhead could not be estimated because single GPU calculations were not feasible due to memory limitations.
Memory usage analysis, panels (e) and (f), and Table 8 shows that duplicated memory remains consistently around 2 GB per GPU, which is 4 times higher than for the CPU case. The distributed memory is very close to the one obtained in the CPU case, see Table 6. On Leonardo Booster, per each GPU there are 64 GB of memory, and for the cases of medium and big matrices, this imposes a minimal amount of GPUs needed to perform the simulations. In comparison with the CPU case, on GPUs the SLEPc solver shows similar or even higher memory consumption than the ELPA solver. This is due to the large parallel overhead memory that we are presently investigating. For the GPU implementation, the parallel overhead memory is of a similar order of magnitude to that observed in the CPU case. Specifically, for SLEPc, the parallel overhead memory is large and positive, indicating the allocation of additional variables for parallel runs. Conversely, for ELPA, the parallel overhead is negative, suggesting more efficient parallelization schemes.